OpenStudy (anonymous):

multiply (3x-4)(6x+7)

OpenStudy (anonymous):

(3x-4)(6x+7)

OpenStudy (anonymous):

18x^2-3x-28 x(18x-3)-28 1/8(12x-1)^2-225/8 x=-7/6 x=4/3

OpenStudy (anonymous):

Distribute the terms (3x-4) * (6x+7) = (3x * 6x) + (3x * 7) + (-4 * 6x) + (-4 * 7) = 18x^2 + 21x -24x -28 = 18x^2 -3x -28

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem. i think kumar set the equation equal to 0, which you don't need unless it tells you to do that. just make sure you know what you're doing for when a test comes